Monday Morning Litho status report 11/29

Gary Sosa gsosa at stanford.edu
Mon Nov 29 07:38:57 PST 2010


Good Morning All.....

    Here is the Monday morning Litho status report.

   Karl Suss 2- No BSA image on right side. Will investigate on morning 
rounds

   Headway- User reports vacuum problem. Will test out on morning rounds

   Laurel- User reports that some chucks do not work. Possible O-ring 
problem. Will check out on morning rounds

   SVG Coat2- User reports dispens earm not moving. Will check out on 
morning rounds

   EVG Bonder- User reports a Graphite plate issue. Will investigate.

      Please check back in Coral for latest tool updates and current 
equipment status.


          Thanks...  The Litho Team

*/Photolithography Area Status/*

	

	

*/Monday, November 29, 2010/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

Mario- Out today

*/Exposure Tools/**//*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*UP*

	

	

EVAlign

	

*UP*

	

	

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*UP*

	

	

Karl Suss 2

	

*PROBLEM*

	

No "Right" BSA image

	

3

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*UP*

	

	

Headway

	

*PROBLEM*

	

Vacuum problem

	

2

Laurel

	

*PROBLEM*

	

Some chucks don't work

	

5

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Up*

	

	

3617 Process

	

*Up*

	

	

220-7 Process

	

*Problem*

	

Optimizing process

	

6

SVGCoat2

	

3612 Process

	

*Problem*

	

Dispense arm issue

	

1

955-0.7 Process

	

*Up*

	

	

220-3 Process

	

*Up*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*UP*

	

	

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*PROBLEM*

	

Temperature Drift. Controller on order

	

7

Blue M 150 C

	

*PROBLEM*

	

Temperature Drift. Controller on order

	

8

Fusion

	

*DOWN*

	

Hotplate experiments

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*UP*

	

	

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*UP*

	

	

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*UP*

	

	

#3 Olympus/Litho

	

*UP*

	

	

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*PROBLEM*

	

Graphite plate issue

	

4

KSBonder

	

*UP*

	

	

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

	

	

	

	

	

	


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20101129/b4f1de4f/attachment.html>


More information about the litho mailing list